OpenStudy (anonymous):

What is the easiest way to find the intersection point of: (x)^3 = cube root of {x+3} ?

OpenStudy (anonymous):

Btw, they are inverses of each other.

OpenStudy (anonymous):

WAIT!

OpenStudy (anonymous):

One minute, I messed up the beginning equation

OpenStudy (anonymous):

its cube on left side and cuberoot on other side?

OpenStudy (anonymous):

its actually suppose to be: (x^3 - 1) = cube root of (x + 1)

OpenStudy (anonymous):

DISREGARD THE FIRST POST; It is suppose to be: (x^3 - 1) = cube root of (x + 1)

OpenStudy (anonymous):

What is the easiest way of finding the intersection pt?

hero (hero):

Cube both sides

OpenStudy (anonymous):

Ok then..

OpenStudy (anonymous):

it is not that easy.. Try trial and error method by plugging in different values for x

OpenStudy (anonymous):

So it can't be solved out easily through algebra?

hero (hero):

graph it

hero (hero):

graph x^3 - 1 then graph cube root(x+3)

hero (hero):

Then find the intersection point

OpenStudy (anonymous):

Ok so I am suppose to guesstimate here?

OpenStudy (anonymous):

Yes,,, you can estimate.. or use calculator to graph two equations.. one the left side and the other the right side// Answer would be 1.32472

OpenStudy (anonymous):

You guys have contradicting answers, but its OK. I will just plot it on WolframAlpha and see what I get. Thanks in advance.

hero (hero):

I graphed it and found the intersection point

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28x^3+-+1%29+%3D+ \sqrt[3]{x%2B1}

hero (hero):

Shucks. Had warrior posted it correctly the first time....

hero (hero):

That explains why we had conflicting answers

hero (hero):

Intersection point is (1.32, 1.32)